# Preserving Expert-Level Privacy in Offline Reinforcement Learning

### v1

@@ 6 EXPERIMENTS
Ensuring reasonable levels of privacy requires significant amounts of hyperparameter tuning and a large number of data points, and typically comes with a non-trivial performance impact [^3] In supervised learning, for example, the accuracy on ImageNet training can go from $>75\%$ to $\sim 47.9\%$ in the private case (Kurakin et al.,, 2022). Intuitively, with the demand for expert-level privacy in our setting, we risk severe degradation of utility since each “record” (each expert’s data) forms a much larger fraction of the overall dataset. It is thus imperative that we create a dataset which consists of diverse data collected from a large number of different experts. We describe our data generation procedure to ensure this below.
6.1 Dataset Generation
@@
We train $3000$ experts each on the Cartpole, Acrobot, LunarLander and HIV Treatment environments. Cartpole, Acrobot and LunarLander are from the Gymnasium package (Towers et al.,, 2023). The HIV Treatment simulator is based on the implementation by Geramifard et al., (2015) of the model described in Ernst et al., (2006). Experts are trained on variations of the default environment, created by modifying the values of key parameters (e.g. gravity in LunarLander environment). The trained experts are then used to generate demonstrations on a default setting of the environment, collectively forming the aggregated dataset for offline RL. This training methodology ensures that the experts show varied (often sub-optimal) behaviours on the default environment. As shown in Figure 2, the experts’ return-histograms indicate that the mix of experts we create shows significant diversity in the policies learnt. Further details regarding the environment settings used to train the expert policies are avilable in Appendix B. We acknowledge that due to Assumption 1, we do not consider continuous control domains here and focus on discrete action benchmarks.
6.2 Training Setup
@@
While our algorithm allows us to use any off-the-shelf gradient-based offline RL algorithm, we use Conservative Q-Learning (CQL) (Kumar et al.,, 2020) algorithm for our experimental work. More specifically, we use the DQN version of CQL, since we are dealing with discrete action spaces. To represent the Q-value function, we use a neural network with 2 hidden layers of 256 units each. During training, we perform a grid search to find the optimal set of hyper-parameters (learning rate, sampling probability $p$, DP-SGD noise) for each environment. The batch size is fixed to 256 for Acrobot, LunarLander and Cartpole, and to 64 for HIV Treatment.
@@
The maximum trajectory length is fixed to 200 for Acrobot and CartPole in the collected demonstrations. While evaluating the final learned policy, the maximum episode length is again set as 200 for Acrobot, we increase this to 1000 for Cartpole for increased difficulty in an otherwise “simple” environment. We allow all DP-SGD training runs to progress until the privacy budget is used up completely and report statistics of episode returns averaged over $10$ evaluation runs at the end of training.
We keep the privacy budget constant to $\varepsilon=10$, and $\delta=1/m$, where $m$ is the number of experts.
6.3 Baselines
We compare our method against DP-SGD adapted for expert-level privacy. A brief discussion of the changes we made are presented in Section 4. A full discussion, along with analysis, is presented in the appendix for reference. We keep the underlying RL algorithm as CQL for fair comparison. Since there are no prior methods for the expert-level privacy setting, we do not have other readily available baselines.
We also compare against a non-private baseline, which ignores privacy concerns and trains over the whole dataset using vanilla CQL. This may be taken as an upper-bound on the performance that we hope to achieve in the presence of privacy concerns.
@@ 6.4 Results
(a) LunarLander
(b) Acrobot
(c) CartPole
(d) HIV Treatment
Figure 3: Our method performs favourably compared to expert-level DP-SGD on multiple test environments. The average episodic return over $10$ evaluation runs at the end of training are reported along with respective confidence intervals. For HIV Treatment, the return values are normalised by $61e$ 6.
(a) LunarLander
(b) Acrobot
(c) CartPole
Figure 4: Performance of our method and DP-SGD for different values of $\varepsilon$. We see consistent improvements of our method over DP-SGD in the higher $\varepsilon$ regions.
@@
Our experiments show that our algorithm which leverages expert consensus to select stable trajectory prefixes prior to training CQL, shows superior performance to training CQL by merely using an adapted expert-level DP-SGD algorithm. Our method is able to recover much of the performance loss due to DP-SGD which is evident by our results being very close to the non-private training results (Figure 3).
@@
We also provide the values of sampling probability that led to the best performance for each environment in Table 1. Interestingly, for Acrobot and HIV Treatment environments, just $D_{stable}^{\Pi}$ was enough to get the optimal performance and we did not need to pay any privacy cost of DP-SGD during model training.
@@
Table 1: Best choices for probability of sampling from $\mathcal{D}^{\Pi}_{stable}$ ie. $1-p$ for Algorithm 3 on test environments
Environment Sampling probability
LunarLander 0.2
Acrobot 1.0
CartPole 0.1
HIV Treatment 1.0
Figure 4 compares the performance of our method with DP-SGD across different values of $\varepsilon$. All the results are averaged over 3 runs and $95\%$ confidence intervals are also reported. Note that our method degenerates to DP-SGD if we take $\varepsilon_{1}=0,\varepsilon_{2}=\varepsilon$ for any given $\varepsilon$, which is what we do in the lower $\varepsilon$-region ($\varepsilon\leq 5$ for Lunar Lander and CartPole, and $\varepsilon\leq 2.5$ for Acrobot) where the privacy budget is not enough to get sufficient stable data for training.
For our method, for higher values of $\varepsilon$, we take $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4}$ for LunarLander and CartPole. For Acrobot, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0$, that is, we only train on $D_{stable}^{\Pi}$ for Acrobot. Similarly, we split $\delta$ as $\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$.
7 CONCLUSION AND FUTURE WORK
@@
In this work, we initiate the study of offline RL with expert-level privacy. We describe the novel challenges of the setting, and provide theoretically sound algorithms which are evaluated in proof-of-concept evaluation. In future work, we would like to relax Assumptions 1-3, which limit the applicability of our results. We expect replacing the queries to the experts with behaviour cloned versions to be less challenging than eliminating the other two assumptions. Understanding the utility implications of Algorithm 3, despite the influence of our privatization scheme on the training distribution is another important avenue for future work. Finally, while our empirical evaluation convincingly establishes the promise of this approach, we leave a larger evaluation across diverse and larger-scale RL benchmarks for future studies.
@@ Appendix A Theorem Proofs
$=\frac{1}{m}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi)\enspace.$ (10)
@@
$\Pr(E_{\tau_{k}}|\Pi^{\prime})=\frac{1}{m^{\prime}}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi^{\prime})\enspace.$ (11)
@@
$|count_{\tau_{k}}(\Pi)-count_{\tau_{k}}(\Pi^{\prime})|\leq 1\quad\text{and}\quad|m-m^{\prime}|=1\enspace.$ (12)
@@
$=\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)\enspace.$ (15)
@@
$\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\enspace.$ (17)
@@
$\frac{\Pr(E_{\tau_{k}}|\Pi)}{\Pr(E_{\tau_{k}}|\Pi^{\prime})}<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\leq e^{\varepsilon^{\prime}}\enspace,$ (18)
@@
[Using Laplace Tail Bound]
Accounting for the randomness of $v$, from (21), (22) and the pdf of the Laplace distribution, we have,
[…]
@@
This gives the statement of the lemma.
@@
$\Pr(\mathcal{A}(\Pi)=\tau_{k})<\Pr(E_{\tau_{k}}|\Pi)\delta^{\prime}$
@@
[…]
@@
afor-$\forall a$, Using $count_{\tau}(\Pi)<c_{min}<\theta$, (since $\theta=c_{min}/p_{min}$), Lemma 5.2, and (26),
@@
$count_{\tau_{k}}(\Pi)<\frac{c_{min}}{p_{min}}=\theta$ (32)
The output of $\mathcal{A}_{2}$ is the output of a sequence of calls to $\mathcal{A}_{PQ}$ (Algorithm 1). The probability that $\mathcal{A}_{2}$ outputs the sequence $r$ can be upper-bounded by the probability of getting $\top$ on the $k^{th}$ call to $\mathcal{A}_{PQ}$. Using this fact along with (32) and Lemma 5.2, we can say,
@@
If $count_{\tau_{L}}(\Pi)<c_{min}$, then using $count_{\tau}(\Pi)<c_{min}<c_{min}/p_{min}=\theta$, (44) and Lemma 5.2, we get,
@@
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}$ $\leq\sum_{\tau\in\mathcal{T}}\Pr(E_{\tau}|\Pi)\delta^{\prime}$ (51)
$=L\delta^{\prime}=\frac{\delta_{1}}{2T}$ (52)
This proves 48, and hence the statement of the theorem.
@@ B.1 Dataset Generation
Lunar Lander: Among the various variable which govern the dynamics of this environment, we chose to modify gravity (default = 9.8), wind power (default = 0.0) and turbulence (default = 0.0) to train different experts. We modified gravity over 10 equally spaced values from 9.0 to 11.0, wind power over 10 equally spaced values between 0.0 to 5.0, and similarly turbulence over 10 equally spaced values from 0.0 to 0.5, to generate 1000 environment settings. For each of these environment settings, we trained agents using the PPO algorithm for $61e$ 6 updates along with hyperparameter tuning of learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). For each environment setting, we chose the top 3 agents that performed the best on the corresponding environment, to form our 3000 experts.
@@
CartPole: For CartPole, we chose to vary gravity, the magnitude of the force exerted when the cart is pushed and the mass of the cart. The default values of gravity, force magnitude and mass of the cart are 9.8, 10.0 and 1.0 respectively. We varied each control over 10 equally spaced values over a fixed range. For gravity, we fixed this range as (8.75, 11.0), for force magnitude as (9.0, 11.25) and for mass of the cart as (0.8, 1.25). Similar to above, for each of these 1000 environment settings, we trained agents using the PPO algorithm for $61e$ 6 updates while tuning the learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). We chose the top 3 agents from each setting to form our 3000 experts.
@@
Acrobot: For this environment, we varied the lengths of the 2 connected links and their masses. The default values of all these parameters is 1.0. We assign the same length to both the links in each setting. We vary this length, the mass of link 1 and the mass of link 2 over 10 equally spaced values across fixed ranges for each. We fix these ranges as (0.8, 1.2) for the length of the links, (0.9, 1.1) for the mass of link 1 and (0.9, 1.1) for the mass of link 2. The combination of these gives us 1000 environment settings. Again, we trained agents using PPO for $61e$ 6 updates over multiple learning rates and batch sizes and picked the top 3 agents from each setting to form our 3000 experts. Each agent is a neural network with hidden-layers of sizes (512, 256, 64).
@@
HIV Treatment: In this domain, the 6-dimensional continuous state represents the concentrations of various cells and viruses. The actions correspond to the dosages of two drugs. The reward is a function of the health outcomes of the patient, and it also penalises the side-effects due to the quantity of drugs prescribed. Hence, the default reward function has 2 penalty terms, a penalty of 20000 on the dosage of drug 1, and a penalty of 2000 on the dosage of drug 2 (Ernst et al., (2006)). In the default setting of the environment, we assume that the treatment at each steps is applied for 20 days and that there are 50 time steps in an episode. To train experts, we vary the penalty on drug 1 over 50 equally spaced values from 18750 to 21250, and the penalty on drug 2 over 20 equally spaced values from 1850 to 2150. Another variation that we introduce is whether the efficacy of the administered drug is noisy or not. So we either add no noise to the drug efficacy, or we add gaussian noise with standard deviation 0.01. All these modifications give us 2000 different environment settings. We trained agents using PPO for $62e$ 6 updates on each of the environment settings and over multiple learning rates and batch sizes. Each agent is represented by a neural network with a single hidden layer of size 512. We then picked the top agents from each of the 2000 environment settings to create experts. We picked 1000 more top-performing agents from the 2nd best agents of all the environment settings to get 3000 experts in total.

### v2

@@ 6 EXPERIMENTS
Ensuring reasonable privacy requires significant hyperparameter tuning and a large number of data points, often impacting performance adversely [^3] In supervised learning, for example, the accuracy on ImageNet training can go from $75\%$ to $\sim 47.9\%$ in the private case (kurakin2022training), under example-level privacy.. Intuitively, with the demand for expert-level privacy in our setting, we risk severe degradation of utility since each “record” (each expert’s data) forms a much larger fraction of the overall dataset. We recover most of this performance loss, significantly outperforming DP-SGD.
@@
Dataset Generation: We train $3000$ experts each on the Cartpole, Acrobot, LunarLander and HIV Treatment environments. Cartpole, Acrobot and LunarLander are from the Gymnasium package (farama2023gym). The HIV Treatment simulator is based on the implementation by geramifard2015rlpy of the model described in ernst2006clinical. Experts are trained on variations of the default environment, created by modifying the values of key parameters (e.g. gravity in LunarLander environment). The trained experts are then used to generate demonstrations on a default setting of the environment, collectively forming the aggregated dataset for offline RL. This training methodology ensures that the experts show varied (often sub-optimal) behaviours on the default environment. As shown in Figure 2, the experts’ return-histograms indicate that the mix of experts we create shows significant diversity in the policies learnt. Further details regarding the environment settings used to train the expert policies are available in Appendix B.
Figure 2: Expert return histograms, with Kernel Density Estimation, for experts trained on heterogeneous (left to right) LunarLander, Acrobot, CartPole and HIV Treatment environments. We use experts with wide ranges of performance on the test environment. Return values for HIV treatment normalised by $10^{6}$.
@@
Training Setup: Our algorithm allows us to use any off-the-shelf gradient-based offline RL algorithm, and we demonstrate this by experimenting with Conservative Q-Learning (CQL) (kumar2020conservative) and Batch Constrained Deep Q-Learning (BCQ) (fujimoto2019off). More specifically, we use the DQN (Deep Q-Network) version of CQL, and the discrete-action version of BCQ (fujimoto2019benchmarking), since we operate under the assumption of discrete action spaces. To represent each of the Q-value functions in both the algorithms and the generative model in BCQ, we use a neural network with 2 hidden layers of 256 units each. During training, we perform a grid search to find the optimal set of hyper-parameters (learning rate, batch size, sampling probability $p$, DP-SGD noise) for each environment; see Appendix B.
@@
The maximum trajectory length is fixed to 200 for Acrobot and CartPole in the collected demonstrations. While evaluating the final learned policy, the maximum episode length is again set as 200 for Acrobot, we increase this to 1000 for Cartpole for increased difficulty in an otherwise “simple” environment. We allow all DP-SGD training runs to progress until the privacy budget is used up completely.
Baselines: We compare our method against DP-SGD adapted for expert-level privacy. A brief discussion of the changes we made are presented in Section 4. A full discussion, along with analysis, is presented in Appendix B. We keep the underlying RL algorithm same for fair comparison. In the absence of prior methods for the expert-level privacy setting, we limit our empirical studies to this naive baseline. We also train a non-private policy over the whole dataset using vanilla CQL/BCQ (ignoring privacy concerns). This serves as an empirical upper-bound on the performance achievable under privacy constraints. In our experiments, we report episodic returns, normalized to lie between 0 (random policy) and 1 (optimal policy) averaged over 10 evaluation runs at the end of training, as a fraction of the return of the non-private baseline.
@@ 6.1 Results
Figure 3: Performance of our method and DP-SGD for different values of $\varepsilon$ with $\delta=1/m$, where $m$ is the number of experts. We report the episodic return, normalized between 0 (random policy) and 11 (optimal policy) averaged over 1010 evaluation runs (with $95\%$ confidence intervals) at the end of training, as a fraction of the non-private baseline. Our method consistently outperforms DP-SGD, especially in the high $\varepsilon$ regions.
@@
We observe empirically that our algorithm, which leverages expert consensus to select stable trajectory prefixes prior to offline RL training, shows superior performance to merely using an adapted expert-level DP-SGD algorithm. Figure 3 compares the performance of our method with DP-SGD across different values of $\varepsilon$. All the results are averaged over 3 runs and $95\%$ confidence intervals are also reported. Our method is able to recover much of the performance loss due to DP-SGD, especially in the high $\varepsilon$-regions.
We experiment with $\varepsilon<5$ too, but for such low $\varepsilon$, Algorithm 2 yields very little stable data, which does not meaningfully influence training. Thus, the model primarily learns from the unstable data. Thus, setting $\varepsilon_{1}=0$ and $\varepsilon_{2}=\varepsilon$ is optimal as we minimize the noise injected by the DP-SGD training. This is in contrast to the observations for higher $\varepsilon$ values where we take $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4}$ for LunarLander and CartPole. For Acrobot and HIVTreatment, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0$, that is, we only train on $D_{stable}^{\Pi}$. Similarly, we split $\delta$ as $\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$. These results for low $\varepsilon$ are presented in Table 2.
@@
We also note the optimal settings of sampling probabilities in Table 1. Interestingly, for Acrobot (with CQL and BCQ) and HIV Treatment (with CQL) environments, just $D_{stable}^{\Pi}$ suffices to get the optimal performance in high $\varepsilon$-region and we did not need to pay any privacy cost of DP-SGD during model training. We also report the variation of performance with various mixes of the two datasets (by varying $p$) in Table 2.
Environment CQL BCQ $\varepsilon=5$ $\varepsilon=7.5$ $\varepsilon=10$ $\varepsilon=5$ $\varepsilon=7.5$ $\varepsilon=10$ LunarLander 1.0 0.8 0.8 1.0 0.5 0.5 Acrobot 0.0 0.0 0.0 0.0 0.0 0.0 CartPole 1.0 0.9 0.8 1.0 0.5 0.5 HIV Treatment 1.0 0.0 0.0 1.0 0.9 0.9
@@
Table 1: Best choices for probability of sampling from $\mathcal{D}^{\Pi}_{unstable}$ ie. $p$ for Algorithm 3 on test environments
Environment CQL BCQ $\bm{p=0.0}$ $\bm{p=0.8}$ $\bm{p=1.0}$ $\bm{p=0.0}$ $\bm{p=0.5}$ $\bm{p=1.0}$ LunarLander inf 0.75 0.67 inf 0.57 0.54 CartPole 0.67 1.03 0.92 0.31 0.99 0.67
Environment $\bm{\varepsilon}$ $\bm{\varepsilon_{1}=\frac{3}{4}\varepsilon}$ $\bm{\varepsilon_{1}=0}$ Acrobot (BCQ) 2.5 0.74 0.95 Cartpole (BCQ) 2.5 0.28 0.58 Cartpole (BCQ) 5.0 0.56 0.60
Table 2: We perform ablation studies with low, intermediate and high sampling probabilities $p$ from the unstable dataset $D^{\Pi}_{unst}$, at $\varepsilon=10$. inf (infeasible) denotes that the stable dataset $D^{\Pi}_{stable}$ was empty, and training was infeasible. We also vary the allocation of $\varepsilon$ in the low-privacy regime ($\varepsilon\leq 5$) and report utility.
We also evaluate the effect of varying the number of experts on the performance of our algorithm relative to DP-SGD, in figure 4, on the Acrobot environment. As $m$ increases, the likelihood that $count_{\tau}(\Pi)$ exceeds the noisy threshold $\hat{\theta}$ also increases, and with it the number and length of trajectory prefixes in the stable dataset $D_{stable}^{\Pi}$. Thus, performance scales with the number of experts. Conversely, for very low values of $m$, the stable dataset is empty and our algorithm defaults towards the baseline.
Figure 4: Performance of our method and DP-SGD for $m=1000,2000$ and 30003000 experts, on the Acrobot environment with $\varepsilon=10.0$ and CQL as the underlying offline RL algorithm. We observe increasing improvements of our method over DP-SGD as $m$ increases.
Note on Implementation: In the implementation of both our method and DP-SGD, we create batches by by shuffling datasets of individual experts and randomly interleaving them such that each batch has at most one data point from each expert. This differs slightly from the poisson sampling based accounting in our privacy analysis. This gap between analysis and implementation in our work corresponds to the gap between privacy amplification by subsampling and shuffling. A promising line of work (erlingsson2019shuffling; feldman2021hiding) aims to derive comparable theoretical amplification guarantees via shuffling. Empirically, however, the gap in utility is expected to be minimal (ponomareva23dpfy).
We use the PLD accountant from the dp-accounting library to account for the privacy loss in DP-SGD training.
7 CONCLUSION
@@
In this work, we initiate the study of offline RL with expert-level privacy. We describe the novel challenges of the setting, and provide theoretically sound algorithms which are evaluated in proof-of-concept evaluation. In future work, we would like to relax Assumptions 3.4-3.6, by replacing the queries to the experts with behaviour cloned versions and extending our trajectory prefix selection paradigm to continuous action spaces (for e.g using kernel density estimation). Understanding the utility implications of Algorithm 3, despite the influence of our privatization scheme on the training distribution is another important avenue for future work. Further, privacy auditing in the reinforcement learning setting is another important open research problem orthogonal to our work. Finally, while our empirical evaluation convincingly establishes the promise of this approach, we leave a larger evaluation across diverse and larger-scale RL benchmarks for future studies.
Acknowledgements
The authors thank Krishna Pillutla and Balaraman Ravindran for fruitful discussions and helpful comments. We also thank the TMLR reviewers for their helpful feedback. This research was supported in part by the Post-Baccalaureate Fellowship at the Centre for Responsible AI, IIT Madras.
@@ Appendix A Theorem Proofs
$=\frac{1}{m}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi)\kern 5.0pt.$ (10)
@@
$\Pr(E_{\tau_{k}}|\Pi^{\prime})=\frac{1}{m^{\prime}}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi^{\prime})\kern 5.0pt.$ (11)
@@
$|count_{\tau_{k}}(\Pi)-count_{\tau_{k}}(\Pi^{\prime})|\leq 1\quad\text{and}\quad|m-m^{\prime}|=1\kern 5.0pt.$ (12)
@@
$=\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)\kern 5.0pt.$ (15)
@@
$\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\kern 5.0pt.$ (17)
@@
$\frac{\Pr(E_{\tau_{k}}|\Pi)}{\Pr(E_{\tau_{k}}|\Pi^{\prime})}<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\leq e^{\varepsilon^{\prime}}\kern 5.0pt,$ (18)
@@
Using the Laplace tail bound, we account for the randomness of $v$ from equation 21, equation 22 and the probability density function of the Laplace distribution, we have,
$\Pr\left[Lap(4/\varepsilon^{\prime})>(4/\varepsilon^{\prime})\log(1/\delta^{\prime})+Lap(2/\varepsilon^{\prime})\right]$ $\leq\int_{-\infty}^{\infty}\left(\frac{\delta^{\prime}}{2}e^{-v(\varepsilon^{\prime}/4)}\times\frac{\varepsilon^{\prime}}{4}e^{-|v|(\varepsilon^{\prime}/2)}\right)dv$ (23)
@@
This gives the statement of the lemma. ∎
@@
$\Pr(\mathcal{A}(\Pi)=\tau_{k})<\Pr(E_{\tau_{k}}|\Pi)\cdot\delta^{\prime}$
@@
$\mathcal{A}_{2}$ runs the Sparse Vector Algorithm using the the query sequence $Q^{\tau_{k}\cdot a}$ to output a sequence of the form $\top\cdots\top\bot$. Here, each $\top$ or $\bot$ is an output of $\mathcal{A}_{PQ}$ (Algorithm 1). The probability that $\mathcal{A}_{2}$ outputs the sequence $r$ can be upper-bounded by the probability of getting $\top$ on the $k^{th}$ call to $\mathcal{A}_{PQ}$. Formally,
@@
$\forall a$, Using $count_{\tau}(\Pi)<c_{min}<\theta$, (since $\theta=c_{min}/p_{min}$), Lemma 5.2, and equation 26,
@@
$\Rightarrow\,\,$ $count_{\tau_{k}}(\Pi)<\frac{c_{min}}{p_{min}}=\theta$ (32)
The output of $\mathcal{A}_{2}$ is the output of a sequence of calls to $\mathcal{A}_{PQ}$ (Algorithm 1). The probability that $\mathcal{A}_{2}$ outputs the sequence $r$ can be upper-bounded by the probability of getting $\top$ on the $k^{th}$ call to $\mathcal{A}_{PQ}$. Using this fact along with equation 32 and Lemma 5.2, we can say,
@@
If $count_{\tau_{L}}(\Pi)<c_{min}$, then using $count_{\tau}(\Pi)<c_{min}<c_{min}/p_{min}=\theta$, equation 44 and Lemma 5.2, we get,
@@
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}$ $\leq\sum_{\tau\in\mathcal{T}}\Pr(E_{\tau}|\Pi)\delta^{\prime}=L\delta^{\prime}=\frac{\delta_{1}}{2T}$ (51)
This proves Eqn 48, and hence the statement of the theorem.
@@ B.1 Dataset Generation
LunarLander: Among the various variable which govern the dynamics of this environment, we chose to modify gravity (default = 9.8), wind power (default = 0.0) and turbulence (default = 0.0) to train different experts. We modified gravity over 10 equally spaced values from 9.0 to 11.0, wind power over 10 equally spaced values between 0.0 to 5.0, and similarly turbulence over 10 equally spaced values from 0.0 to 0.5, to generate 1000 environment settings. For each of these environment settings, we trained agents using the PPO algorithm for $1e6$ updates along with hyperparameter tuning of learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). For each environment setting, we chose the top 3 agents that performed the best on the corresponding environment, to form our 3000 experts.
@@
CartPole: For CartPole, we chose to vary gravity, the magnitude of the force exerted when the cart is pushed and the mass of the cart. The default values of gravity, force magnitude and mass of the cart are 9.8, 10.0 and 1.0 respectively. We varied each control over 10 equally spaced values over a fixed range. For gravity, we fixed this range as (8.75, 11.0), for force magnitude as (9.0, 11.25) and for mass of the cart as (0.8, 1.25). Similar to above, for each of these 1000 environment settings, we trained agents using the PPO algorithm for $1e6$ updates while tuning the learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). We chose the top 3 agents from each setting to form our 3000 experts.
@@
Acrobot: For this environment, we varied the lengths of the 2 connected links and their masses. The default values of all these parameters is 1.0. We assign the same length to both the links in each setting. We vary this length, the mass of link 1 and the mass of link 2 over 10 equally spaced values across fixed ranges for each. We fix these ranges as (0.8, 1.2) for the length of the links, (0.9, 1.1) for the mass of link 1 and (0.9, 1.1) for the mass of link 2. The combination of these gives us 1000 environment settings. Again, we trained agents using PPO for $1e6$ updates over multiple learning rates and batch sizes and picked the top 3 agents from each setting to form our 3000 experts. Each agent is a neural network with hidden-layers of sizes (512, 256, 64).
@@
HIV Treatment: In this domain, the 6-dimensional continuous state represents the concentrations of various cells and viruses. The actions correspond to the dosages of two drugs. The reward is a function of the health outcomes of the patient, and it also penalises the side-effects due to the quantity of drugs prescribed. Hence, the default reward function has 2 penalty terms, a penalty of 20000 on the dosage of drug 1, and a penalty of 2000 on the dosage of drug 2 (ernst2006clinical). In the default setting of the environment, we assume that the treatment at each steps is applied for 20 days and that there are 50 time steps in an episode. To train experts, we vary the penalty on drug 1 over 50 equally spaced values from 18750 to 21250, and the penalty on drug 2 over 20 equally spaced values from 1850 to 2150. Another variation that we introduce is whether the efficacy of the administered drug is noisy or not. So we either add no noise to the drug efficacy, or we add gaussian noise with standard deviation 0.01. All these modifications give us 2000 different environment settings. We trained agents using PPO for $2e6$ updates on each of the environment settings and over multiple learning rates and batch sizes. Each agent is represented by a neural network with a single hidden layer of size 512. We then picked the top agents from each of the 2000 environment settings to create experts. We picked 1000 more top-performing agents from the 2nd best agents of all the environment settings to get 3000 experts in total.